How do I solve 2x+3y=18 and x=2y-5 its one problem and I want to use the substitution method.thanks
Found 2 solutions by scott8148, Theo:
Answer by scott8148(6628)

substitute the equivalent value of x from the 2nd equation into the 1st equation

2(2y - 5) + 3y = 18 ___ 4y - 10 + 3y = 18 ___ 7y = 28

solve for y, then substitute back to find x

Answer by Theo(13342)

2x + 3y = 18
x = 2y-5
substitute 2y-5 for x in the first equation to get:
2(2y-5) + 3y = 18
simplify by removing parentheses to get:
4y - 10 + 3y = 18
combine like terms to get:
7y - 10 = 18
add 10 to both sides to get:
7y = 28
divide both sides by 7 to get:
y = 4
substitute 4 for y in the second equation of x = 2y-5 to get:
x = 2*4-5 which becomes:
x = 8-5 which becomes:
x = 3
substitute 3 for x and 4 for y in the first equation of 2x + 3y = 18 to get:
2*3 + 3*4 = 18 which becomes:
6 + 12 = 18 which becomes:
18 = 18
values for x and y are good.
x = 3
y = 4